15. Where does diminishing marginal product begin? Workers Total Product
A. after first worker buite color B. after second worker C. after third worker ingen soro D. after fourth worker or nobod
Answer #1

Answer
marginal product of n th worker =total output of n workers - total output of n-1 workers
MP(n)=TP(n)-TP(n-1)

MP(1)=3-0=3
MP(2)=8-3=5
MP(3)=10-8=2
MP(4)=11-10=1
MP(5)=9-11=-2
MP(6)=6-9=-3

the diminishing marginal product begins as the marginal product decreases and it is decreased after the second worker

Option B
after the second worker